OpenStudy (anonymous):

An automobile travels at 100km/h. a) Find the angular velocity of a tire with radius 36cm. b) Through what angle will the tire turn in 30s at this speed?

OpenStudy (perl):

you there?

OpenStudy (anonymous):

yep

OpenStudy (perl):

ok

OpenStudy (perl):

there is a nice equation relating linear speed and angular speed (or angular velocity). linear velocity = radius * angular velocity (in radians). this comes from the equation s = r *theta, divide both sides by t. s/t = r * theta/t , and you get the equation

OpenStudy (anonymous):

This is a bit confusing still...do you have a full solution? the answer says 77.2rad/min

OpenStudy (perl):

so we are given

OpenStudy (perl):

An automobile travels at 100km/h. a) Find the angular velocity of a tire with radius 36cm. b) Through what angle will the tire turn in 30s at this speed?

OpenStudy (perl):

so using v = r * w , where v = linear speed, r = radius, and w = angular speed in radians/ unit time. we have 100 km/h = 36 cm * w rad/ time .

OpenStudy (perl):

now we have kilometers and centimeters, which don't cancel. but know that 110 km = 110*1000*100 cm

OpenStudy (perl):

ok so based on your solution it looks like they want radians per minute, so we have to do another conversion

OpenStudy (anonymous):

but it says angular velocity, what do you do from there?

OpenStudy (perl):

first of all, you have a mistake. you want angular velocity in radians per second

OpenStudy (anonymous):

but the answer given is 77.2r/min

OpenStudy (perl):